OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 36 inches. Find the width of the rectangle.

OpenStudy (anonymous):

w = x l=x+4 P=2L+2w Solve for x 2(x+4) + 2x=36

OpenStudy (anonymous):

7?